Round your answers to the nearest hundredth.




Triangle QRS has vertices Q(1, 6), R (7, 9), and S (10, 6). What is the approximate perimeter of triangle QRS?
1
Expert's answer
2013-05-16T09:58:34-0400
Let's calculate the lengths to estimate the perimeter of the triangle. We'll use Pythagoras' theorem to do that:

QR = √((1-7)² + (6-9)²)& = √(36 + 9) = √45 ≈ 6.71
RS = √((7-10)² + (9-6)²) = √(9 + 9)& = √18 ≈ 4.24
SQ = √((10-1)² + (6-6)²) = √(81 + 0) = 9

So, the perimeter of the triangle is approximately

P = QR + RS + SQ = 6.71 + 4.24 + 9 = 19.95